# FocalCodec: Low-Bitrate Speech Coding via Focal Modulation Networks

### v1

@@ 3.2 Training
The training process consists of two stages. In the first stage, the compressor, quantizer, and decompressor are jointly trained to reconstruct the encoder continuous representations, ensuring that the tokens retain both semantic and acoustic information from the encoder, which is kept frozen. The training objective includes reconstruction loss and entropy loss. The reconstruction loss is computed as the squared L2 distance between the reconstructed and original encoder features. The entropy loss, defined as in (Yu et al., 2024; Zhao et al., 2024), encourages both confident predictions and uniform code utilization. Note that we omit the commitment loss used in standard VQ, as for BSQ there is no concern of embedding divergence (quantization error is bounded).
In the second stage, the decoder is trained to resynthesize audio from the encoder continuous representations. This approach enables us to perform this stage in parallel with the first, simplifying the training setup. The training objective includes adversarial loss, reconstruction loss, and feature matching loss, as in (Kong et al., 2020). However, following (Zeghidour et al., 2021), we use a hinge loss formulation instead of least squares. The reconstruction loss is computed as the L1 distance between the reconstructed and original log-Mel spectrograms, while the feature matching loss is the mean of the distances between the $l$-th feature maps of the $k$-th subdiscriminator.
This decoupled training approach ensures that both semantic and acoustic information are preserved in the tokens, which is crucial for downstream tasks while maintaining high reconstruction quality. If trained end-to-end without additional constraints on the hidden representations (e.g. distillation loss), the reconstruction loss prioritizes acoustic features, as observed in (Défossez et al., 2023; Kumar et al., 2023).
@@ 4.1 FocalCodec
We train FocalCodec on LibriTTS (Zen et al., 2019), resampled to 16 kHz. We train three variants of the model with a codebook size of 8192 and token rates of 50 Hz, 25 Hz, and 12.5 Hz by adjusting the temporal downsampling factors in the compressor layers to (1, 1, 1), (2, 1, 1), and (2, 2, 1), respectively. These patterns are mirrored in the decompressor layers for upsampling. Information about hyperparameters and training details can be found in Section D.1.
@@ 4.2 Baselines
We compare our models to recent state-of-the-art low-bitrate codecs across acoustic, semantic, and hybrid categories. Since the paper focuses on low-bitrate codecs, when multiple quantizers are available, we configure them to achieve a bitrate below 1.50 kbps, ensuring a fair comparison. For acoustic codecs, we compare against EnCodec (Défossez et al., 2023), DAC (Kumar et al., 2023), WavTokenizer (Ji et al., 2024), and BigCodec (Xin et al., 2024). Among these, BigCodec is the current state-of-the-art for low-bitrate speech reconstruction quality (Wu et al., 2024a). We use the official checkpoints for these models. We do not include the recent TS3-Codec (Wu et al., 2024a), which matches BigCodec performance at an even lower bitrate, as it is not publicly available. However, we contacted the authors to request reconstructed samples for comparison. Additional results related to TS3-Codec can be found in Section F.2.
For semantic codecs, we adopt the approach introduced in (Wang et al., 2024), which quantizes layer-6 representations from WavLM-large using k-means clustering with 512 centroids. These representations are fed into a Conformer (Gulati et al., 2020) encoder to reconstruct continuous representations, followed by a HiFi-GAN decoder. This baseline, referred to as WavLM6-KM, provides a direct comparison between our codec and another model leveraging WavLM layer-6 features but differing in design and training methodology. Since the code and checkpoints for WavLM6-KM are not publicly available, we reimplemented the model using a subset of LibriSpeech (Panayotov et al., 2015). Note that we do not include additional baselines from this category, as semantic codecs typically underperform in terms of reconstruction quality (Parker et al., 2024) or require much higher bitrates to be competitive in this regard (Mousavi et al., 2024a). Furthermore, most hybrid codecs are already built on top of semantic representations. Therefore, we prioritize the hybrid category, to which our codec also belongs. For hybrid codecs, we compare against SpeechTokenizer (Zhang et al., 2024), SemantiCodec (Liu et al., 2024), Mimi (Défossez et al., 2024), and Stable Codec (Parker et al., 2024), using their official checkpoints. The configurations and details of each model are summarized in Table 1. Multiply-accumulate operations per second (MACs) are measured using ptflops [^3] https://pypi.org/project/ptflops/0.7.4/. Additional information about the baselines is provided in Appendix C.
Table 1: Compared codecs.
Codec Bitrate (kbps) Sample Rate (kHz) Token Rate (Hz) Codebooks Code Size Params (M) MACs (G)
@@
Stable Codec 0.70 16 25.0 2 $\times$ 15625 6 95 37
@@ 4.3 Speech Resynthesis
We evaluate FocalCodec on speech resynthesis, considering both English and multilingual speech. For English speech, we use LibriSpeech (Panayotov et al., 2015) test-clean. For multilingual speech, following (Xin et al., 2024), we randomly select 100 utterances from each of the 7 foreign languages in Multilingual LibriSpeech (Pratap et al., 2020) (Dutch, French, German, Italian, Polish, Portuguese, and Spanish), resulting in a total of 700 utterances. We also consider the more realistic scenario of speech contaminated with environmental noise. For this, we use the test splits of VoiceBank (Valentini-Botinhao et al., 2016) and the more challenging Libri1Mix, which is constructed by mixing clean utterances from the first speaker of LibriMix (Cosentino et al., 2020) with noise from WHAM! (Wichern et al., 2019).
@@
Table 2: Clean speech resynthesis.
@@
Codec UTMOS $\uparrow$ dWER $\downarrow$ Sim $\uparrow$ Code Usage $\uparrow$ Norm Entropy $\uparrow$ RTF $\uparrow$
LibriSpeech test-clean
@@
EnCodec 1.58 8.08 93.8 93.4 82.1 109
DAC 1.29 20.04 89.2 100.0 91.7 89
WavLM6-KM 3.75 6.20 90.0 26.4 95.4 85
SpeechTokenizer 2.28 5.14 91.6 95.9 97.0 63
SemantiCodec 2.91 8.97 96.0 75.9 94.4 0.62
Mimi 3.29 5.73 96.0 95.6 91.8 137
WavTokenizer 3.78 11.55 95.4 100.0 96.7 181
BigCodec 4.11 2.55 98.5 100.0 98.6 22
Stable Codec 4.32 4.97 94.7 98.5 94.7 103
FocalCodec@50 4.05 2.18 97.4 100.0 98.9 185
FocalCodec@25 4.14 3.30 96.3 99.8 98.4 195
FocalCodec@12.5 4.22 7.94 93.9 98.2 97.4 208
Multilingual LibriSpeech 700
Reference 2.84 0.00 100.0
EnCodec 1.33 29.60 95.5 93.4 79.2 140
DAC 1.24 56.08 89.1 100.0 90.0 97
WavLM6-KM 2.97 44.54 89.5 28.1 0.91 125
SpeechTokenizer 1.55 56.32 92.0 96.1 94.0 74
SemantiCodec 1.87 36.21 97.7 76.4 94.7 0.74
Mimi 2.08 30.96 96.7 95.9 89.0 239
WavTokenizer 2.64 49.73 97.0 97.6 95.6 290
BigCodec 2.86 15.24 99.1 100.0 97.9 24
Stable Codec 3.47 56.99 95.9 92.9 93.8 144
FocalCodec@50 2.96 12.57 98.3 100.0 98.1 269
FocalCodec@25 3.16 19.78 97.3 99.2 97.4 292
FocalCodec@12.5 3.37 54.15 95.2 96.4 96.9 296
Table 3: Noisy speech resynthesis.
Codec DNSMOS $\uparrow$ dWER $\downarrow$ Sim $\uparrow$ Code Usage $\uparrow$ Norm Entropy $\uparrow$ RTF $\uparrow$
VoiceBank test
Reference 3.56 0.00 100.0
EnCodec 2.76 28.16 87.7 77.5 78.1 44
DAC 2.72 63.90 79.8 98.7 88.4 48
WavLM6-KM 3.06 20.67 82.9 24.8 92.3 44
SpeechTokenizer 2.74 34.51 82.2 88.1 88.4 42
SemantiCodec 3.13 31.46 90.6 52.4 92.6 0.28
Mimi 3.01 28.00 87.8 78.6 85.5 47
WavTokenizer 3.09 42.12 89.8 94.8 94.0 63
BigCodec 3.19 20.67 92.3 99.8 96.8 17
Stable Codec 3.33 20.32 88.8 75.7 95.4 39
FocalCodec@50 3.16 8.08 91.3 98.0 96.2 80
FocalCodec@25 3.17 11.75 90.1 89.6 96.0 81
FocalCodec@12.5 3.22 27.97 84.7 77.3 95.5 79
Libri1Mix test
Reference 3.73 0.00 100.0
EnCodec 2.40 55.17 86.3 84.4 78.7 97
DAC 2.40 90.92 76.6 99.1 88.8 91
WavLM6-KM 2.87 36.60 85.9 26.8 95.5 65
SpeechTokenizer 2.58 57.26 82.8 93.5 96.5 63
SemantiCodec 2.67 51.18 89.9 64.7 90.8 91
Mimi 2.65 49.14 89.4 90.8 90.1 104
WavTokenizer 2.53 70.10 86.3 96.4 95.4 165
BigCodec 2.75 53.26 88.3 100.0 98.2 19
Stable Codec 2.91 43.52 90.0 95.8 93.4 68
FocalCodec@50 2.93 27.89 91.6 100.0 98.5 155
FocalCodec@25 2.91 34.27 90.7 99.6 97.9 161
FocalCodec@12.5 2.92 42.59 88.9 97.2 97.2 164
We evaluate the models using objective metrics. To measure naturalness, we employ UTMOS (Saeki et al., 2022) for clean speech and DNSMOS (Reddy et al., 2022) for noisy speech. Note that we do not include signal-level metrics such as SNR, PESQ (Rix et al., 2001), or STOI (Taal et al., 2011), as these metrics do not correlate well with perceived reconstruction quality (Parker et al., 2024; Wang et al., 2024). To evaluate speaker fidelity, we compute the cosine similarity (Sim) between speaker embeddings extracted from the reconstructed audio and the target audio. These embeddings are obtained using WavLM-base (Chen et al., 2022) fine-tuned for speaker verification [^4] https://huggingface.co/microsoft/wavlm-base-sv. To assess intelligibility, we compute the differential word error rate (dWER(Wang et al., 2021), which measures the difference in word error rate between the reconstructed and target audio, using transcriptions from Whisper small [^5] https://huggingface.co/openai/whisper-small (Radford et al., 2023). To ensure fairness in evaluation, we do not use more powerful ASR models (e.g. Whisper large-v3), as these models can correct pronunciation mistakes and are more robust to noise, potentially hiding flaws in the reconstruction. We also report code usage, i.e. the ratio of unique tokens used to the codebook size (averaged over codebooks for multi-codebook models), and normalized entropy (Cover & Thomas, 2006; Parker et al., 2024), where higher values indicate more uniform codebook usage. For inference speed, we measure the real-time factor (RTF), i.e. the ratio of the reconstructed audio duration to the processing time. An RTF greater than 1 indicates faster-than-real-time performance, measured on an NVIDIA V100 GPU with 32 GB of memory.
Discussion
Results are presented in Table 2 and Table 3. FocalCodec shows strong performance across both clean and noisy speech resynthesis tasks. On clean speech, FocalCodec@50 achieves the best trade-off of quality, intelligibility, and efficiency. Notably, FocalCodec is the best in terms of dWER, surpassing BigCodec, which is currently state-of-the-art. It also generalizes well to multilingual speech, obtaining the lowest dWER and high Sim. Note that FocalCodec, WavLM6-KM, SpeechTokenizer, BigCodec and Stable Codec were trained exclusively on English speech. In noisy speech resynthesis, FocalCodec@50 again excels, achieving the lowest dWER by a large margin on both VoiceBank and Libri1Mix, while maintaining high speaker similarity. Meanwhile, FocalCodec@25 and FocalCodec@12.5 exhibit some degradation in dWER and speaker similarity, particularly in multilingual settings, due to their significantly lower bitrates. Nevertheless, despite operating at just 0.16 kbps, FocalCodec@12.5 remains competitive with several baselines that use much higher bitrates (e.g. EnCodec). Finally, the high code usage and normalized entropy across all FocalCodec variants indicate efficient token utilization, contributing to their strong overall performance. Additional results on reconstruction quality, including subjective evaluations, can be found in Section F.1.
Table 4: One-shot voice conversion.
Codec UTMOS $\uparrow$ dWER $\downarrow$ Sim $\uparrow$ RTF $\uparrow$
VCTK
Reference 4.09 0.00 100.0
EnCodec 1.24 86.52 72.2 57
DAC 1.25 104.00 67.2 60
WavLM6-KM 2.90 26.68 92.4 57
SpeechTokenizer 1.49 20.32 81.2 33
SemantiCodec 2.02 106.00 72.8 0.60
Mimi 2.40 110.00 89.7 71
WavTokenizer 3.13 43.15 73.4 89
BigCodec 1.31 99.96 68.9 13
Stable Codec 3.76 27.63 71.1 65
FocalCodec@50 3.38 21.27 92.2 116
FocalCodec@25 3.40 23.59 92.6 118
FocalCodec@12.5 3.43 29.93 92.6 117
4.4 Voice Conversion
We conduct one-shot voice conversion experiments to verify that FocalCodec can effectively disentangle speaker information from content despite its single-codebook design. This task involves converting speech from a source speaker to an arbitrary target speaker using reference speech from the target speaker. For single-codebook baselines, including FocalCodec, we use $k$-nearest neighbors search in the codec feature space, as in (Baas et al., 2023). Specifically, we replace each frame in the reconstructed feature sequence (right before the decoder) with the average of the $k=\text{4}$ closest matches in terms of cosine distance from continuous features extracted from the reference. For multi-codebook baselines, instead, we follow the procedure in (Zhang et al., 2024). The source and reference speech are tokenized, and the first codebook tokens from the source are concatenated with the second-to-last codebook tokens from the reference. The resulting sequence is then forwarded to the decoder. If sequence lengths differ, the reference is truncated or circularly padded as needed. Effective disentanglement of content and speaker information between first and subsequent codebooks is expected to yield fair voice conversion performance. We conduct voice conversion experiments on VCTK (Yamagishi et al., 2017), which includes parallel utterances from different speakers. To create the test set, we randomly select an utterance from a source speaker, the corresponding utterance from a target speaker, and an utterance with different content from the same target speaker to act as the reference. Among available reference utterances, we select the longest to minimize padding issues. We repeat this process for each speaker, for each of the $\sim\text{24}$ parallel utterances, resulting in a dataset with 2521 samples. To evaluate performance, we use UTMOS, dWER, Sim, and RTF as defined in Section 4.3.
Discussion
As reported in Table 4, FocalCodec achieves the highest speaker similarity while maintaining good intelligibility, confirming its suitability for voice conversion tasks. This is particularly impressive, especially compared to other hybrid codecs like SpeechTokenizer and Mimi, which are explicitly optimized to disentangle semantic information in the first codebook and acoustic information in the following. Despite this, FocalCodec outperforms these models, excelling in both speaker identity preservation and intelligibility, striking a remarkable balance of quality, efficiency, and speaker similarity. WavLM6-KM ranks as the second-best performing model, which is expected since it shares the same encoder as FocalCodec. In contrast, acoustic codecs struggle with this task, as they do not separate speaker and content information.
Table 5: Evaluation on downstream tasks.
Codec Discriminative Tasks Generative Tasks
ASR SI SER SE SS TTS
Reference 3.56 0.00 100.0 3.77 0.00 100.0 4.09 0.00 100.0
EnCodec 27.89 3.00 47.00 3.11 37.10 85.9 3.11 78.51 87.3 1.69 74.07 79.1
DAC 35.89 3.27 45.90 3.03 67.65 81.7 2.76 106.00 83.3 1.36 61.11 84.1
WavLM6-KM 19.04 22.30 42.90 3.52 22.85 83.6 3.49 76.91 85.0 3.71 48.51 88.2
SpeechTokenizer 14.97 2.73 41.50 3.21 29.82 85.9 3.13 83.99 87.3 2.63 47.81 88.3
SemantiCodec 41.42 15.90 51.60 3.59 102.00 83.3 3.59 123.00 84.4 2.72 59.85 90.8
Mimi 22.98 5.43 44.70 3.30 53.98 84.6 3.41 93.23 88.1 3.05 39.50 93.3
WavTokenizer 35.62 2.44 49.80 3.41 51.75 88.6 3.54 105.00 86.4 3.65 59.22 89.6
BigCodec 26.41 2.34 47.50 3.52 26.68 93.2 3.54 89.24 89.4 3.24 63.83 87.8
Stable Codec 16.85 16.50 46.54 3.55 35.57 82.8 3.61 103.00 78.2 2.86 56.97 84.3
FocalCodec@50 17.63 4.48 45.60 3.47 10.93 91.4 3.71 73.87 89.0 4.05 39.58 92.9
FocalCodec@25 21.12 6.07 46.80 3.49 14.74 90.0 3.69 99.96 85.4 4.12 30.28 91.4
FocalCodec@12.5 33.24 11.69 46.30 3.58 36.98 86.9 3.57 116.00 80.8 4.16 29.91 91.5

### v2

@@ 3.2 Training
The training process consists of two stages. In the first stage, the compressor, quantizer, and decompressor are jointly trained to reconstruct the encoder continuous representations, ensuring that the tokens retain both semantic and acoustic information from the encoder, which is kept frozen. The training objective includes reconstruction loss and entropy loss. The reconstruction loss is computed as the squared L2 distance between the reconstructed and original encoder features. The entropy loss, defined as in yu2024lfq ; zhao2024bsq , encourages both confident predictions and uniform code utilization. Note that we omit the commitment loss used in standard VQ, as for BSQ there is no concern of embedding divergence (quantization error is bounded).
In the second stage, the decoder is trained to resynthesize audio from the encoder continuous representations. The training objective includes adversarial loss, reconstruction loss, and feature matching loss, as in kong2020hifigan . However, following zeghidour2021soundstream , we use a hinge loss formulation instead of least squares. The reconstruction loss is computed as the L1 distance between the reconstructed and original log-Mel spectrograms, while the feature matching loss is the mean of the distances between the $l$-th feature maps of the $k$-th subdiscriminator.
This design allows the second stage to run in parallel with the first, simplifying the training setup. At inference, the same decoder operates on dequantized features produced by the compressor-quantizer-decompressor pipeline. Because the decompressor is trained to reconstruct the original continuous features from the discrete codes, these dequantized features closely approximate the originals. As a result, the decoder maintains strong performance even when using dequantized features as input, without requiring any additional fine-tuning.
This decoupled training approach ensures that both semantic and acoustic information are preserved in the tokens, which is crucial for downstream tasks while maintaining high reconstruction quality. If trained end-to-end without additional constraints on the hidden representations (e.g. distillation loss), the reconstruction loss prioritizes acoustic features, as observed in defossez2023encodec ; kumar2023dac .
@@ 4.1 FocalCodec
We train FocalCodec on LibriTTS zen2019libritts , resampled to 16 kHz. We train three variants of the model with a codebook size of 8192 and token rates of 50 Hz, 25 Hz, and 12.5 Hz by adjusting the temporal downsampling factors in the compressor layers to (1, 1, 1), (2, 1, 1), and (2, 2, 1), respectively. These patterns are mirrored in the decompressor layers for upsampling. Information about hyperparameters and training details can be found in Section˜E.1.
@@ 4.2 Baselines
Table 1: Codecs considered in our experimental analysis.
Codec
Bitrate
Sample
Rate
Token
Rate
Codebooks
Code
Size
Params
MACs
@@
Stable Codec 0.70 16 25.0 2 $\times$ 15625 6 950 37
@@
We compare our models to recent state-of-the-art low-bitrate codecs across acoustic, semantic, and hybrid categories. Since this paper focuses on low-bitrate codecs, when multiple quantizers are available, we configure them to achieve a bitrate below 1.50 kbps, ensuring a fair comparison. For acoustic codecs, we compare against EnCodec defossez2023encodec , DAC [^2] Note that we use the 16 kHz checkpoint, whereas the original results in kumar2023dac use the 24 kHz checkpoint. kumar2023dac , WavTokenizer ji2024wavtokenizer , and BigCodec xin2024bigcodec . Among these, BigCodec is the current state-of-the-art for low-bitrate speech reconstruction quality wu2024ts3codec . We use the official checkpoints for these models. We do not include the recent TS3-Codec wu2024ts3codec , which matches BigCodec performance at an even lower bitrate, as it is not publicly available. However, we contacted the authors to request reconstructed samples for comparison. Additional results related to TS3-Codec can be found in Section˜G.1.
For semantic codecs, we adopt the approach introduced in wang2024selm , which quantizes layer-6 representations from WavLM-large using k-means clustering with 512 centroids. These representations are fed into a Conformer gulati2020conformer encoder to reconstruct continuous representations, followed by a HiFi-GAN decoder. This baseline, referred to as WavLM6-KM, provides a direct comparison between our codec and another model leveraging WavLM layer-6 features but differing in design and training methodology. Since the code and checkpoints for WavLM6-KM are not publicly available, we reimplemented the model using a subset of LibriSpeech panayotov2015librispeech . Note that we do not include additional baselines from this category, as semantic codecs typically underperform in terms of reconstruction quality parker2024scaling or require much higher bitrates to be competitive in this regard mousavi2024dasb . Furthermore, most hybrid codecs are already built on top of semantic representations. Therefore, we prioritize the hybrid category, to which our codec also belongs. For hybrid codecs, we compare against SpeechTokenizer zhang2024speechtokenizer , SemantiCodec liu2024semanticodec , Mimi defossez2024moshi , and Stable Codec parker2024scaling , using their official checkpoints. The configurations and details of each model are summarized in Table˜1. Multiply-accumulate operations per second (MACs) are measured using ptflops [^3] https://pypi.org/project/ptflops/0.7.4/. Additional information about the baselines is provided in Appendix˜D.
@@ 4.3 Speech Resynthesis
We evaluate FocalCodec on speech resynthesis, considering both English and multilingual speech. For English speech, we use LibriSpeech panayotov2015librispeech test-clean. For multilingual speech, following xin2024bigcodec , we randomly select 100 utterances from each of the 7 foreign languages in Multilingual LibriSpeech pratap2020mls (Dutch, French, German, Italian, Polish, Portuguese, and Spanish), resulting in a total of 700 utterances [^4] https://zenodo.org/records/14791114. We also consider the more realistic scenario of speech contaminated with environmental noise. For this, we use the test splits of VoiceBank valentinibotinhao2016voicebank and the more challenging Libri1Mix, which is constructed by mixing clean utterances from the first speaker of LibriMix cosentino2020librimix with noise from WHAM! wichern2019wham .
We evaluate the models using objective metrics. To measure naturalness, we employ UTMOS saeki2022utmos for clean speech and DNSMOS reddy2022dnsmos for noisy speech. Note that we do not include signal-level metrics such as SNR, PESQ rix2001pesq , or STOI taal2011stoi , as these metrics do not correlate well with perceived reconstruction quality parker2024scaling ; wang2024selm . To evaluate speaker fidelity, we compute the cosine similarity (Sim) between speaker embeddings extracted from the reconstructed audio and the target audio. These embeddings are obtained using WavLM-base-SV [^5] https://huggingface.co/microsoft/wavlm-base-sv chen2022wavlm . To assess intelligibility, we compute the differential word error rate (dWERwang2021dwer , which measures the difference in word error rate between the reconstructed and target audio, using transcriptions from Whisper small [^6] https://huggingface.co/openai/whisper-small radford2022robust . To ensure fairness in evaluation, we do not use more powerful ASR models (e.g. Whisper large-v3), as these models can correct pronunciation mistakes and are more robust to noise, potentially hiding flaws in the reconstruction. We also report code usage, i.e. the ratio of unique tokens used to the codebook size (averaged over codebooks for multi-codebook models), and normalized entropy thomas2006elements ; parker2024scaling , where higher values indicate more uniform codebook usage. For inference speed, we measure the real-time factor (RTF), i.e. the ratio of the reconstructed audio duration to the processing time. An RTF greater than 1 indicates faster-than-real-time performance, measured on an NVIDIA V100 GPU with 32 GB of memory.
Results are presented in Table˜2. FocalCodec shows strong performance across both clean and noisy speech resynthesis tasks. On clean speech, FocalCodec@50 achieves the best trade-off of quality, intelligibility, and efficiency. Notably, FocalCodec is the best in terms of dWER, surpassing BigCodec, which is currently state-of-the-art. It also generalizes well to multilingual speech, obtaining the lowest dWER and high Sim. Note that FocalCodec, WavLM6-KM, SpeechTokenizer, BigCodec and Stable Codec were trained exclusively on English speech. In noisy speech resynthesis, FocalCodec@50 again excels, achieving the lowest dWER by a large margin on both VoiceBank and Libri1Mix, while maintaining high speaker similarity. Meanwhile, FocalCodec@25 and FocalCodec@12.5 exhibit some degradation in dWER and speaker similarity, particularly in multilingual settings, due to their significantly lower bitrates. Nevertheless, despite operating at just 0.16 kbps, FocalCodec@12.5 remains competitive with several baselines that use much higher bitrates (e.g. EnCodec). It is also worth noting that FocalCodec’s UTMOS tends to increase at lower bitrates, likely due to the stronger smoothing effect introduced by downsampling. However, UTMOS tends to saturate and may not fully capture perceptual quality saeki2022utmos . dWER and Sim are therefore essential to provide a more comprehensive evaluation. Finally, the high code usage and normalized entropy across all FocalCodec variants indicate efficient token utilization, contributing to their strong overall performance. Additional results on reconstruction quality, including subjective evaluations, streamability and Mel-spectrogram analysis, can be found in Sections˜G.2, G.3 and G.4.
@@
Table 2: Speech resynthesis performance across clean and noisy datasets.
@@
Codec Bitrate (kbps) $\downarrow$ UTMOS $\uparrow$ dWER $\downarrow$ Sim $\uparrow$ Code Usage $\uparrow$ Norm. Entropy $\uparrow$ RTF $\uparrow$ DNSMOS $\uparrow$ dWER $\downarrow$ Sim $\uparrow$ Code Usage $\uparrow$ Norm. Entropy $\uparrow$ RTF $\uparrow$
Clean – LibriSpeech test-clean Noisy – VoiceBank
Reference 4.09 0.00 100.0 3.56 0.00 100.0
EnCodec 1.50 1.58 8.08 93.8 93.4 82.1 109 2.76 28.16 87.7 77.5 78.1 44
DAC 1.00 1.29 20.04 89.2 100.0 91.7 89 2.72 63.90 79.8 98.7 88.4 48
WavLM6-KM 0.45 3.75 6.20 90.0 26.4 95.4 85 3.06 20.67 82.9 24.8 92.3 44
SpeechTokenizer 1.00 2.28 5.14 91.6 95.9 97.0 63 2.74 34.51 82.2 88.1 88.4 42
SemantiCodec 0.65 2.91 8.97 96.0 75.9 94.4 0.62 3.13 31.46 90.6 52.4 92.6 0.28
Mimi 0.69 3.29 5.73 96.0 95.6 91.8 137 3.01 28.00 87.8 78.6 85.5 47
WavTokenizer 0.48 3.78 11.55 95.4 100.0 96.7 181 3.09 42.12 89.8 94.8 94.0 63
BigCodec 1.04 4.11 2.55 98.5 100.0 98.6 22 3.19 20.67 92.3 99.8 96.8 17
Stable Codec 0.70 4.32 4.97 94.7 98.5 94.7 103 3.33 20.32 88.8 75.7 95.4 39
FocalCodec @50 0.65 4.05 2.18 97.4 100.0 98.9 185 3.16 8.08 91.3 98.0 96.2 80
FocalCodec @25 0.33 4.14 3.30 96.3 99.8 98.4 195 3.17 11.75 90.1 89.6 96.0 81
FocalCodec @12.5 0.16 4.22 7.94 93.9 98.2 97.4 208 3.22 27.97 84.7 77.3 95.5 79
Clean – Multilingual LibriSpeech 700 Noisy – Libri1Mix
Reference 2.84 0.00 100.0 3.73 0.00 100.0
EnCodec 1.50 1.33 29.60 95.5 93.4 79.2 140 2.40 55.17 86.3 84.4 78.7 97
DAC 1.00 1.24 56.08 89.1 100.0 90.0 97 2.40 90.92 76.6 99.1 88.8 91
WavLM6-KM 0.45 2.97 44.54 89.5 28.1 0.91 125 2.87 36.60 85.9 26.8 95.5 65
SpeechTokenizer 1.00 1.55 56.32 92.0 96.1 94.0 74 2.58 57.26 82.8 93.5 96.5 63
SemantiCodec 0.65 1.87 36.21 97.7 76.4 94.7 0.74 2.67 51.18 89.9 64.7 90.8 91
Mimi 0.69 2.08 30.96 96.7 95.9 89.0 239 2.65 49.14 89.4 90.8 90.1 104
WavTokenizer 0.48 2.64 49.73 97.0 97.6 95.6 290 2.53 70.10 86.3 96.4 95.4 165
BigCodec 1.04 2.86 15.24 99.1 100.0 97.9 24 2.75 53.26 88.3 100.0 98.2 19
Stable Codec 0.70 3.47 56.99 95.9 92.9 93.8 144 2.91 43.52 90.0 95.8 93.4 68
FocalCodec @50 0.65 2.96 12.57 98.3 100.0 98.1 269 2.93 27.89 91.6 100.0 98.5 155
FocalCodec @25 0.33 3.16 19.78 97.3 99.2 97.4 292 2.91 34.27 90.7 99.6 97.9 161
FocalCodec @12.5 0.16 3.37 54.15 95.2 96.4 96.9 296 2.92 42.59 88.9 97.2 97.2 164
4.4 Voice Conversion
We conduct one-shot voice conversion experiments to verify that FocalCodec can effectively disentangle speaker information from content despite its single-codebook design. This task involves
Table 3: One-shot voice conversion on VCTK veaux2017cstr .
Codec
Bitrate
$\downarrow$
UTMOS $\uparrow$
dWER $\downarrow$
Sim $\uparrow$
RTF $\uparrow$
@@ 4.4 Voice Conversion
EnCodec 1.50 1.24 86.52 72.2 57
DAC 1.00 1.25 104.00 67.2 60
WavLM6-KM 0.45 2.90 26.68 92.4 57
SpeechTokenizer 1.00 1.49 20.32 81.2 33
SemantiCodec 0.65 2.02 106.00 72.8 0.60
Mimi 0.69 2.40 110.00 89.7 71
WavTokenizer 0.48 3.13 43.15 73.4 89
BigCodec 1.04 1.31 99.96 68.9 13
Stable Codec 0.70 3.76 27.63 71.1 65
FocalCodec @50 0.65 3.38 21.27 92.2 116
FocalCodec @25 0.33 3.40 23.59 92.6 118
FocalCodec @12.5 0.16 3.43 29.93 92.6 117
converting speech from a source speaker to an arbitrary target speaker using reference speech from the target speaker. For single-codebook baselines, including FocalCodec, we use $k$-nearest neighbors search in the codec feature space, as in baas2023knnvc . Specifically, we replace each frame in the reconstructed feature sequence (right before the decoder) with the average of the $k$ = 4 closest matches in terms of cosine distance from continuous features extracted from the reference. For multi-codebook baselines, instead, we follow the procedure in zhang2024speechtokenizer . The source and reference speech are tokenized, and the first codebook tokens from the source are concatenated with the second-to-last codebook tokens from the reference. The resulting sequence is then forwarded to the decoder. If sequence lengths differ, the reference is truncated or circularly padded as needed. Effective disentanglement of content and speaker information between first and subsequent codebooks is expected to yield fair voice conversion performance. We conduct voice conversion experiments on VCTK veaux2017cstr , which includes parallel utterances from different speakers. To create the test set, we randomly select an utterance from a source speaker, the corresponding utterance from a target speaker, and an utterance with different content from the same target speaker to act as the reference. Among available reference utterances, we select the longest to minimize padding issues. We repeat this process for each speaker, for each of the $\sim$24 parallel utterances, resulting in a dataset with 2521 samples. To evaluate performance, we use UTMOS, dWER, Sim, and RTF as defined in Section˜4.3.
As reported in Table˜3, FocalCodec achieves the highest speaker similarity while maintaining good intelligibility, confirming its suitability for voice conversion tasks. This is particularly impressive, especially compared to other hybrid codecs like SpeechTokenizer and Mimi, which are explicitly optimized to disentangle semantic information in the first codebook and acoustic information in the following. Despite this, FocalCodec outperforms these models, excelling in both speaker identity preservation and intelligibility, striking a remarkable balance of quality, efficiency, and speaker similarity. WavLM6-KM ranks as the second-best performing model, which is expected since it shares the same encoder as FocalCodec. In contrast, acoustic codecs struggle with this task, as they do not separate speaker and content information.